Our systems administrators have identified an issue in the RAID for iis517. The server is currently under a raid rebuild. We will keep you updated on the status as often as possible.
- 2/06/2013 9:52 AM - The operating system has be reinstalled and our admins are currently restoring the disk data.
- 2/06/2013 1:32 PM -We are finishing installing server applications. ETA currently 16 hours for total server recreation.
- 2/07/2013 7:30 AM -Our administrators are preparing to install the control panel. Current ETA is an additional 10 hours.
- 2/07/2013 5:08 PM - Our administrators are working on seting up the IPs and then into the recreation of IIS site structure. Current ETA is 10 hours.
- 2/08/2013 02:56 AM - Recreation slower than expected. Current ETA is 12 hours.
- 2/08/2013 01:36 PM - Recreation is still ongoing. Current ETA is 10-24 hours.
- 2/09/2013 1:446 PM - Recreation of user accounts via control panel has resulted in errors that are currently being investigated. Once resolved, account recreation can resume. ETA is unfortunately increased to 48 hours.
- 2/09/2013 11:00 PM - Recreation is still ongoing. Our current ETA is 30 hours.
- 2/10/2013 9:52 AM - Physical recreation is on the way. 400 domains have been recreated. Around 900 domains left. Once all domains are recreated our admins will start working on recovering permissions. ETA 32 hours domains recreation + recovering permissions. We will have an ETA on recovering permissions once it starts.
- 2/10/2013 7:41 PM - Recreation of all the domains is still underway. The recreation of domains is set for about another 22 hours then there will be some time to correct all permissions on the server. Estimated ETA for everything is 40 hours.
- 2/11/2013 12:56 PM - Recreation of all the domains is still underway. The recreation of domains is set for about another 22 hours then there will be some time to correct all permissions on the server. Estimated ETA for everything is still around 40 hours.
- 2/11/2013 3:15 PM - Currently we are working on copying all user content back to there respective home directories for all customers on the server. ETA is 30 hours for all data to be restored completely.
- 2/12/2013 7:33 AM - We are restoring content from a local copy. We have 60 accounts done and 120 more to go and then our administrators will have to restore permissions. ETA for account restoration is 16 ho0urs plus the time it will take to restore permissions.
- 2/12/2013 2:20 PM - Content has been restored. We are still installing other control panel features and restoring permissions. ETA 30 hr.
- 2/13/2013 06:16 AM Restoring permissions on accounts. ETA for this stage is 40 hours.
- 2/13/2013 12:42 PM 7 out of 155 accounts have had permissions restored. After this there are 2 more steps in the server recreation before it will be back online. Current ETA is 35 hours.
- 2/13/2013 8:12 PM Current ETA is down to 30 hours. Our administrators are working diligently to restore permissions.
- 2/14/2013 05:59 AM Permission restoration continues. Current ETA is about 24 hours.
- 2/14/2013 11:24 AM Permission restoration continues. 75 accounts have been restored with 80 still remaining. ETA 24 hours.
- 2/14/2013 3:50 PM Permission restoration continues. 90 out of 155 accounts have had permissions restored. ETA 20 hours.
- 2/14/2013 9:10 PM Permission restoration continues. 130 out of 155 accounts have had permissions restored. ETA 16 hours. Once this is complete complete server will be started and site functionality verified. If all is well and no further troubleshooting is needed, accounts will be made accessible at that time.
- 2/15/2013 12:25 AM Permission restoration complete. Sites should be available, however they may experience slowness or interruptions. We are troubleshooting and verifying all server details. Current ETA is about 12 hours.
- 2/15/2013 10:04 AM - Troubleshooting and verification of all server details continue. Our administrators are focusing on bringing sites online. Secondary services like FTP will be brought online as soon as possible. Current ETA 10 hours.
- 2/15/2013 3:10 PM - Troubleshooting and verification continue. Current concern is an issue with network connectivity stability. In order to isolate the issue, some services (including web service) will need to be toggled. This will impact site availability. Current ETA 6 hours.
- 2/15/2013 8:10 PM - All services are up and running. Some fixes need to be aplied to PHP but these should be completed before the end of the previous ETA. Current ETA less than 2 hours.
- 2/16/2013 7:15 AM - The server is completely back online. There may be some issues with SSL configuration within individual accounts. If you are having any issues please call, chat or put in a ticket with us 24/7 and we would be happy to get that resolved.
The server is back online. If you are experiencing any issues please call, chat or put in a ticket with us.